# Traversable wormholes inside anisotropic magnetized neutron stars: physical properties and potential observational imprints

### v1

@@ IV Non-magnetized configuration of WH+NS systems
By using the expression of $A$, $C$, and $V$ shown by Eqs. (26)-(28), we are left with
@@
which is the conservation law for anisotropic fluid. For non-magnetized configuration, we have
@@
Note that in this work we use the polytropic equation of state (EOS), which reads [31, 38, 79, 80, 81]
@@
In this paper, we adopt the same parameters as those used in Ref. [38], namely $K=100~\mathrm{km}^{2}$ and $n=1$. Therefore, the EOS reduces to $p_{r}=K\rho^{2}$. For the matter-energy density profile, we also follow Ref. [38] and employ a Tolman–VII-like profile, i.e.
@@
Generally, the total mass of the WH+NS systems is given by [38]
@@
Since $\rho_{\text{tot}}=0$, the total mass is only taken from the ADM mass contribution, i.e. Eq. (44). It is also important to highlight that for the isotropic case where $h=1$, the calculations recover to the one presented in Ref. [38].
@@
Both approaches have also been employed in Ref. [38], which serves as a methodological predecessor to the procedure adopted in this work. In the present analysis, we consider $h=1,;1.5,;2$.
For all considered values of $h$ within Approach 1, as we can see in Fig. 1(a), the mass of the WH+NS systems can exceed that of the secondary compact object in GW190814, indicating that the resulting configurations are exceptionally massive. Similar mass enhancement has already been reported for isotropic systems in Ref. [38]. However, our inclusion of anisotropy allows the formation of even more massive configurations, accompanied by a larger surface radius.
@@
Fig. 2(a) shows the comparison between the radius of the WH+NS system and the corresponding wormhole throat radius obtained using Approach 1. The results indicate that the throat radii for different values of the anisotropy parameter $h$ almost coincide, implying that anisotropy does not significantly affect the size of the wormhole throat. Its influence is instead mainly reflected in the accessible range of the throat radius, where configurations with larger $h$ extend to higher mass values.
The wormhole throat radii obtained using Approach 2, shown in Fig. 2(b), exhibit a trend similar to that found in Approach 1. In particular, the throat radius curves again nearly overlap for different values of $h$, indicating a comparable insensitivity of the throat size to anisotropy. The qualitative effect of anisotropy on the throat radius range is also the same as in Approach 1. The primary difference between the two approaches appears in the overall mass and radius scales: Approach 1 yields systematically larger masses and radii, consistent with the corresponding MR relations obtained under that approach. This suggests that the wormhole throat size is primarily governed by the global structure of the configuration rather than by pressure anisotropy.
As mentioned earlier, the WH+NS system can reach masses that may even exceed the mass of the secondary compact object of GW190814. This fact implies that the system can become very massive and, consequently, highly compact, as illustrated in Fig. 3. The compactness $\mathcal{C}=M/R_{s}$ can even enter the region typically associated with ultracompact objects. The range of $\mathcal{C}$ for ultracompact object is $0.33<\mathcal{C}<0.44$ [82, 83]. Recently, research on ultracompact objects has attracted significant interest (see, e.g., Refs. [82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94]). Those objects possess radii smaller than $3M$, i.e., inside the photon-sphere radius. In such regimes, gravitational-wave echoes may occur, and the corresponding echo time can be computed [96], thereby providing a potential observational imprint of such systems. The fact that the system considered in this work can enter the ultracompact region is therefore a notable result, as this regime cannot be achieved by ordinary polytropic NSs.
@@
Some physical quantities can serve as potential observational imprints of WH+NS configurations, one of which is the surface redshift $z$. Mathematically, it is defined as [96]
@@
For ordinary NSs with typical masses $M=1.2\text{--}2.0\,M_{\odot}$ and radii $R_{s}=10\text{--}14$ km, the expected surface redshift falls within the range $z\simeq 0.1\text{--}0.5$. Cottam et al. [97] reported a gravitational redshift of $z=0.35$ from 28 bursts of the low-mass X-ray binary EXO 0748-676, consistent with theoretical expectations for regular NSs.
@@
Another promising potential observational imprint of the WH+NS systems is the echo time. By assuming that the wormhole is connected to an identical object residing in another spacetime—following the setup considered in Ref. [98]—gravitational waves can propagate through the wormhole mouths and become trapped between the two photon spheres. This process gives rise to gravitational echoes (see Fig. 1 of Ref. [98] for a schematic illustration).
@@
Within this framework, the echo time $\tau$ can be expressed as [98]
@@
By combining Eqs. (48) and (49), one obtains $r=3r_{0}/2=3M$, which corresponds to the location of the photon sphere. It is important to emphasize that this result is model dependent. For instance, in alternative compact object configurations such as dark energy stars with phantom fields, the photon sphere radius $R_{p}$ can deviate from the value $3M$. In such scenarios, the upper limit of the integral in Eq. (47) is consequently modified, as demonstrated in Ref. [96].
@@
V Magnetized configuration of WH+NS systems
The EMT for magnetized configuration is shown by Eq. (21). The conservation law gives
@@ V Magnetized configuration of WH+NS systems
$\nu$ $=$ $-2\left(2-\frac{1}{h}\right)\int\frac{d\rho}{\rho+p_{r}+\frac{1}{6\pi}\left(B_{s}+B_{0}\frac{\rho}{\rho_{0}}\right)^{2}}-\frac{1}{12\pi}\int\frac{B\,dB}{\rho_{0}\left(\frac{B-B_{s}}{B_{0}}\right)+K\left[\rho_{0}\left(\frac{B-B_{s}}{B_{0}}\right)\right]^{2}+\frac{B^{2}}{6\pi}}.$ (52)
Now the first and second terms of the right-hand side in Eq. (52) can be defined as $I_{1}$ and $I_{2}$, respectively. We have
$I_{1}=-2\left(2-\frac{1}{h}\right)\int\frac{\rho d\rho}{\alpha\rho^{2}+\beta\rho+\gamma},$ (53)
$I_{2}=-\frac{z}{12}\int\frac{(z\rho+B_{s})d\rho}{\alpha\rho^{2}+\beta\rho+\gamma},$ (54)
where
$\alpha=K+\frac{B_{0}^{2}}{6\pi\rho_{0}^{2}},\qquad\beta=1+\frac{B_{0}B_{s}}{3\pi\rho_{0}},\qquad\gamma=\frac{B_{s}^{2}}{6\pi},\qquad z=\frac{B_{0}}{\rho_{0}}.$ (55)
By defining $D=\beta^{2}-4\alpha\gamma$, the integral in Eq. (53) will give
$\int\frac{\rho\,d\rho}{\alpha\rho^{2}+\beta\rho+\gamma}$ $=$ $\frac{1}{2\alpha}\ln\left(\alpha\rho^{2}+\beta\rho+\gamma\right)\left\{\begin{array}[]{ll}\displaystyle-\,\frac{\beta}{\alpha}\frac{1}{\sqrt{|D|}}\tan^{-1}\!\left(\frac{2\alpha\rho+\beta}{\sqrt{|D|}}\right),&\text{for }D<0,\\[12.91663pt] \displaystyle-\,\frac{\beta}{2\alpha}\,\frac{1}{\sqrt{D}}\ln\!\left(\frac{2\alpha\rho+\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{2\alpha\rho+\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\right),&\text{for }D>0.\end{array}\right$. (58)
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Considering that we face two possibilities for the second term on the right-hand side of Eq. (58), we decide to perform the calculation using four ranges of surface magnetic field strengths, while keeping the range of $\rho_{0}$ fixed between $1\times 10^{-4}$ and $1\times 10^{4}$ MeV fm-3. The four ranges are summarized in Table 1.
@@
The choice of these values is motivated by the fact that the typical surface magnetic field of neutron stars lies in the range $10^{12}$$10^{15}$ G. Furthermore, according to the simulations reported in Ref. [65], a surface field of $B_{s}=10^{15}$ G corresponds to a core magnetic field of $B_{0}=3\times 10^{18}$ G. Therefore, in our analysis we restrict the order of magnitude of $B_{0}$ to be at most three orders of magnitude larger than $B_{s}$. For the values of $\rho_{0}$ used in this work, our non-magnetized pure NS calculations give $\rho_{0}=3240.46~\text{MeV fm}^{-3}$ for $h=1$, $\rho_{0}=2008.27~\text{MeV fm}^{-3}$ for $h=1.5$, and $\rho_{0}=1459.81~\text{MeV fm}^{-3}$ for $h=2$. We see that all these values remain inside the range $\times 10^{-4}1×1$$\times 10^{4}~\text{MeV fm}^{-3}$. Thus, all values of $\rho_{0}$, $B_{s}$, and $B_{0}$ used throughout this study lie within the parameter ranges described above.
@@
In Fig. 6, the gray region corresponds to the range $0\leq D<10^{-40}$, while the black region represents negative values of $D$. Note that $1\times 10^{-40}$ is the maximum threshold of numerical precision in our computation. We observe that no gray or black regions appear in the figure, indicating that the iscriminant remains positive and always above $1\times 10^{-40}$ throughout all computed regimes. This shows that we can use the logarithmic function for the second term on the right-hand side of Eq. (58), instead of the inverse tangent function. Now Eq. (58) reads
$\int\frac{\rho d\rho}{\alpha\rho^{2}+\beta\rho+\gamma}$ $=$ $\frac{1}{2\alpha}\ln\left(\alpha\rho^{2}+\beta\rho+\gamma\right)-\frac{\beta}{2\alpha}\frac{1}{\sqrt{D}}\ln\left(\frac{2\alpha\rho+\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{2\alpha\rho+\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\right).$ (59)
So we have
$I_{1}=\ln\left[(\alpha\rho^{2}+\beta\rho+\gamma)^{1/2\alpha}\bigg(\frac{2\alpha\rho+\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{2\alpha\rho+\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\bigg)^{-\frac{\beta}{2\alpha\sqrt{\beta^{2}-4\alpha\gamma}}}\right]^{-2K\left(2-\frac{1}{h}\right)}+C_{1}.$ (60)
With the same calculation method, we obtain
$I_{2}=\ln\left[(\alpha\rho^{2}+\beta\rho+\gamma)^{z^{2}/2\alpha}\left(\frac{2\alpha\rho+\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{2\alpha\rho+\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\right)^{\frac{z\left(B_{s}-\frac{z\beta}{2\alpha}\right)}{\sqrt{\beta^{2}-4\alpha\gamma}}}\right]^{-\frac{1}{12\pi}}+C_{2}.$ (61)
The metric function $\nu(r)$ is now given by
$\nu=$ $\>\nu_{c}+\ln\left\{\left[(\alpha\rho^{2}+\beta\rho+\gamma)^{\frac{1}{2\alpha}}\left(\frac{2\alpha\rho+\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{2\alpha\rho+\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\right)^{-\frac{\beta}{2\alpha\sqrt{\beta^{2}-4\alpha\gamma}}}\right]^{-2K\left(2-\frac{1}{h}\right)}\right$.
$\qquad\qquad\left.\times\left[(\alpha\rho^{2}+\beta\rho+\gamma)^{\frac{z^{2}}{2\alpha}}\left(\frac{2\alpha\rho+\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{2\alpha\rho+\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\right)^{\frac{z\left(B_{s}-\frac{z\beta}{2\alpha}\right)}{\sqrt{\beta^{2}-4\alpha\gamma}}}\right]^{-\frac{1}{12\pi}}\right\},$ (62)
where $\nu_{c}=C_{1}+C_{2}$ is the integration constant. The function $e^{2\nu}$ can be easily obtain, i.e.
$e^{2\nu}=$ $\>e^{2\nu_{c}}\left\{\left[(\alpha\rho^{2}+\beta\rho+\gamma)^{\frac{1}{2\alpha}}\left(\frac{2\alpha\rho+\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{2\alpha\rho+\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\right)^{-\frac{\beta}{2\alpha\sqrt{\beta^{2}-4\alpha\gamma}}}\right]^{-4K\left(2-\frac{1}{h}\right)}\right$.
$\qquad\qquad\left.\times\left[(\alpha\rho^{2}+\beta\rho+\gamma)^{\frac{z^{2}}{2\alpha}}\left(\frac{2\alpha\rho+\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{2\alpha\rho+\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\right)^{\frac{z\left(B_{s}-\frac{z\beta}{2\alpha}\right)}{\sqrt{\beta^{2}-4\alpha\gamma}}}\right]^{-\frac{1}{6\pi}}\right\}.$ (63)
At $r=R_{s}$, $\rho=0$ and $e^{2\nu(R_{s})}=1-\frac{r_{0}}{R_{s}}$, so we have
$e^{2\nu_{c}}=\left(1-\frac{r_{0}}{R_{s}}\right)\left(\gamma^{1/\alpha}\right)^{\left[4K\left(2-\frac{1}{h}\right)+\frac{z^{2}}{6\pi}\right]}\left[\left(\frac{\beta-\sqrt{\beta^{2}-4\alpha\gamma}}{\beta+\sqrt{\beta^{2}-4\alpha\gamma}}\right)^{\frac{1}{\sqrt{\beta^{2}-4\alpha\gamma}}}\right]^{\left[\frac{B_{s}z}{6\pi}-\frac{\beta}{2\alpha}\left\{\frac{z^{2}}{6\pi}+4K\left(2-\frac{1}{h}\right)\right\}\right]}.$ (64)
In Appendix A, we show that for the isotropic configuration with the limit $\mathcal{B}\rightarrow 0$, $e^{2\nu}$ is reduces to the expression of $e^{2\nu}$ shown by Eq. (24) in Ref. [38].
The next step is calculating the differential of $e^{\nu}$ with respect to $r$, i.e.
$\left(e^{2\nu}\right)^{\prime}=e^{2\nu}\frac{\rho_{c}(r-r_{0})}{(R_{s}-r_{0})^{2}}\left[\frac{4K\left(2-\frac{1}{h}\right)}{\alpha}\left(\frac{2\alpha\rho+\beta-1}{\alpha\rho^{2}+\beta\rho+\gamma}\right)+\frac{1}{3\pi}\left\{\frac{z}{2\alpha}\frac{(2\alpha\rho+\beta)z+\left(B_{s}-\frac{z\beta}{2\alpha}\right)}{\alpha\rho^{2}+\beta\rho+\gamma}\right\}\right].$ (65)
In Appendix B, we show that for the isotropic configuration with the limit $\mathcal{B}\rightarrow 0$, $\left(e^{2\nu}\right)^{\prime}$ is reduces to the expression of $\left(e^{2\nu}\right)^{\prime}$ shown by Eq. (24) in Ref. [38].
@@
Furthermore, by substituting the expressions of $\alpha$, $\beta$, $\gamma$, and $z$ in Eq. (55) into Eq. (66), we are left with
@@
In this magnetized configurations, we also calculate the surface redshift and the echo time.
Note that in this work, regarding the influence of the magnetic field on the EOS, we follow the approach used in Refs. [71, 99, 100, 101, 102], where the magnetic field contribution to the EOS is taken to be negligible. In particular, Refs. [99, 100] also employed a polytropic EOS in their analysis. The authors of Ref. [71] argued that the dominant magnetic effects arise from the additional stress and pressure terms, while the contribution from magnetization remains subdominant even for extremely strong magnetic fields, especially when such high central fields tend to drive the star toward instability. They used magnetic fields of the order of $10^{18}$ G in their analysis. The neglect of Landau quantization effects on the EOS has also been discussed in Ref. [103].
In the numerical calculations of magnetized WH+NS configurations we consider two representative combinations of magnetic fields: (i) $B_{0}=3\times 10^{15}\,\mathrm{G}$ and $B_{s}=10^{18}\,\mathrm{G}$, and (ii) $B_{0}=3\times 10^{18}\,\mathrm{G}$ and $B_{s}=10^{16}\,\mathrm{G}$. Although the typical surface magnetic field of NSs lies in the range $10^{12}$$10^{15}\,\mathrm{G}$ as mentioned in Sec. I, Ref. [104] reports that the limiting magnetic moment of NSs corresponds to a surface field of order $B_{s}\sim 10^{16}\,\mathrm{G}$. For this reason we also include $B_{s}=10^{16}\,\mathrm{G}$ among the cases examined in our numerical calculations.
We perform the numerical calculations for two magnetic-field configurations: (i) $B_{0}=3\times 10^{18}\,\mathrm{G}$ and $B_{s}=1\times 10^{15}\,\mathrm{G}$, and (ii) $B_{0}=3\times 10^{18}\,\mathrm{G}$ and $B_{s}=1\times 10^{16}\,\mathrm{G}$. Both configurations lie within the parameter ranges shown in Fig. 6. The anisotropy parameter is chosen to be the same as in the non-magnetized case, namely h=1, 1.5,h=1,\ 1.5, and $2$.
@@
Another important point is that observational constraints indicate that the typical radius of neutron stars lies in the range of approximately 1010-$11.5$ km [105], while the canonical neutron-star mass is $1.4\,M_{\odot}$ [49]. Therefore, the very large radii in the range 20209090 km obtained in Fig. 7 are incompatible with current observational evidence, at least at the present stage.

### v3

@@ 4 Non-magnetized configuration of WH+NS systems
By using the expression of $A$, $C$, and $V$ shown by Eqs. (LABEL:A)-(LABEL:V), we are left with
@@
which is the conservation law for anisotropic fluid.
It is worth emphasizing that the conservation equation (44) is not imposed as an independent assumption for the neutron fluid alone. Instead, it emerges naturally from the equations of motion obtained by varying the action with respect to the auxiliary scalar fields. This procedure guarantees the internal consistency of the constrained scalar-gravity system that underlies the WH+NS construction.
Physically, this approach ensures that the presence of auxiliary scalar fields, introduced to generate the effective exotic matter sector required by the wormhole geometry, does not spoil the standard conservation properties of the ordinary matter. Although the scalar fields contribute implicitly through the geometry and the effective EMT, the resulting conservation law involves only the ordinary matter and magnetic field variables. In this sense, the conservation equation should be interpreted as a consistency condition of the full system rather than as an externally imposed constraint.
For non-magnetized configuration, we have
@@
Note that in this work we use the polytropic equation of state (EOS), which reads Dzhunushaliev2012 ; Nojiri2024 ; daSilva2021 ; Godani2024 ; Ray2007
@@
In this paper, we adopt the same parameters as those used in Ref. Nojiri2024 , namely $K=100~\mathrm{km}^{2}$ and $n=1$. Therefore, the EOS reduces to $p_{r}=K\rho^{2}$. For the matter-energy density profile, we also follow Ref. Nojiri2024 and employ a Tolman–VII-like profile, i.e.
@@
Generally, the total mass of the WH+NS systems is given by Nojiri2024
@@
Since $\rho_{\text{tot}}=0$, the total mass is only taken from the ADM mass contribution, i.e. Eq. (54). It is also important to highlight that for the isotropic case where $h=1$, the calculations recover to the one presented in Ref. Nojiri2024 .
@@
Both approaches have also been employed in Ref. Nojiri2024 , which serves as a methodological predecessor to the procedure adopted in this work. In the present analysis, we consider $h=1,;1.5,;2$.
For all considered values of $h$ within Approach 1, as we can see in Fig. 1(a), the mass of the WH+NS systems can exceed that of the secondary compact object in GW190814, indicating that the resulting configurations are exceptionally massive. Similar mass enhancement has already been reported for isotropic systems in Ref. Nojiri2024 . However, our inclusion of anisotropy allows the formation of even more massive configurations, accompanied by a larger surface radius.
@@
Fig. 2(a) shows the comparison between the radius of the WH+NS system and the corresponding wormhole throat radius obtained using Approach 1. We emphasize that the coincidence of the wormhole throat radii observed in Figs. 2(a) and (b) is exact. This is a direct consequence of the geometric relation $r_{0}=2M$, cf. Eq. (44). Since the ADM mass $M$ is used as the horizontal axis in the figures, the functional dependence $r_{0}(M)$ is fixed and identical for all configurations.
The role of the anisotropy parameter $h$ and other matter parameters is not to modify the relation between $r_{0}$ and $M$, but to determine the values of the ADM mass that are realized by physically admissible solutions. Different values of $h$ therefore lead to different mass ranges and, consequently, to different accessible throat radii through the relation $M=r_{0}/2$, while preserving the exact coincidence of the $r_{0}(M)$ curves.
The wormhole throat radii obtained using Approach 2, shown in Fig. 2(b), exhibit a trend similar to that found in Approach 1. The primary difference between the two approaches appears in the overall mass and radius scales: Approach 1 yields systematically larger masses and radii, consistent with the corresponding MR relations obtained under that approach.
As mentioned earlier, the WH+NS system can reach masses that may even exceed the mass of the secondary compact object of GW190814. This fact implies that the system can become very massive and, consequently, highly compact, as illustrated in Fig. 3. The compactness $\mathcal{C}=M/R_{s}$ can even enter the region typically associated with ultracompact objects. The range of $\mathcal{C}$ for ultracompact object is $0.33<\mathcal{C}<0.44$ ByonBurikham ; ByonAnto . Recently, research on ultracompact objects has attracted significant interest (see, e.g., Refs. ByonBurikham ; ByonAnto ; Prasetyo2021 ; Prasetyo2021b ; Fauzi2025 ; Carballo2022 ; Benetti2025 ; Urbano2019 ; Prasetyo2023 ; Hidalgo2021 ; Tasinato2022 ; Kleihaus2020 ; Zhu2020 ). Those objects possess radii smaller than $3M$, i.e., inside the photon-sphere radius. In such regimes, gravitational-wave echoes may occur, and the corresponding echo time can be computed Sakti2021 , thereby providing a potential observational imprint of such systems. The fact that the system considered in this work can enter the ultracompact region is therefore a notable result, as this regime cannot be achieved by ordinary polytropic NSs.
@@
Some physical quantities can serve as potential observational imprints of WH+NS configurations, one of which is the surface redshift $z$. Mathematically, it is defined as Sakti2021
@@
For ordinary NSs with typical masses $M=1.2\text{--}2.0\,M_{\odot}$ and radii $R_{s}=10\text{--}14$ km, the expected surface redshift falls within the range $z\simeq 0.1\text{--}0.5$. Cottam et al. Cottam2002 reported a gravitational redshift of $z=0.35$ from 28 bursts of the low-mass X-ray binary EXO 0748-676, consistent with theoretical expectations for regular NSs.
@@
Another promising potential observational imprint of the WH+NS systems is the echo time. By assuming that the wormhole is connected to an identical object residing in another spacetime—following the setup considered in Ref. Lai2025 —gravitational waves can propagate through the wormhole mouths and become trapped between the two photon spheres. This process gives rise to gravitational echoes (see Fig. 1 of Ref. Lai2025 for a schematic illustration).
@@
Within this framework, the echo time $\tau$ can be expressed as Lai2025
@@
By combining Eqs. (58) and (59), one obtains $r=3r_{0}/2=3M$, which corresponds to the location of the photon sphere. It is important to emphasize that this result is model dependent. For instance, in alternative compact object configurations such as dark energy stars with phantom fields, the photon sphere radius $R_{p}$ can deviate from the value $3M$. In such scenarios, the upper limit of the integral in Eq. (57) is consequently modified, as demonstrated in Ref. Sakti2021 .
@@
5 Magnetized configuration of WH+NS systems
Before discussing the WH+NS configurations, we clarify the role of the parameter $\rho_{0}$ used throughout this work. Here, $\rho_{0}$ denotes the central energy density of the corresponding non-magnetized pure NS configurations and is treated as a scanning parameter.
The range $10^{-4}\leq\rho_{0}\leq 10^{4}\,\mathrm{MeV\,fm^{-3}}$ is chosen to encompass the physically relevant domain of central energy densities typically considered in NS structure calculations, from low-density solutions to very dense cores. Thus, $\rho_{0}$ must exist within this range.
For each value of the anisotropy parameter $h$, the value of $\rho_{0}$ is determined numerically from the corresponding pure neutron-star solutions. Specifically, we solve the standard TOV equations for non-magnetized neutron stars by varying the central energy density $\rho_{c}$ and identify the value of $\rho_{c}$ that yields the maximum-mass configuration.
For the specific values of the anisotropy parameter considered in this work, the maximum-mass pure NS configurations are found at $\rho_{0}=3240.46$ MeV fm-3 for $h=1$, $\rho_{0}=2008.27$ MeV fm-3 for $h=1.5$, and $\rho_{0}=1459.81$ MeV fm-3 for $h=2$, respectively. All these values lie well within the range $10^{-4}\leq\rho_{0}\leq 10^{4}\,\mathrm{MeV\,fm^{-3}}$ adopted in our analysis.
The EMT for magnetized configuration is shown by Eq. (LABEL:EMTmagnetized). The conservation law gives
@@ 5 Magnetized configuration of WH+NS systems
Considering that we face two possibilities for the second term on the right-hand side of Eq. (LABEL:integrallongI1), we decide to perform the calculation using four ranges of surface magnetic field strengths, while keeping the range of $\rho_{0}$ fixed between $1\times 10^{-4}$ and $1\times 10^{4}$ MeV fm-3. The four ranges are summarized in Table 1.
@@
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
The choice of these values is motivated by the fact that the typical surface magnetic field of NSs lies in the range $10^{12}$$10^{15}$ G. Furthermore, according to the simulations reported in Ref. Lopes2020 , a surface field of $B_{s}=10^{15}$ G corresponds to a core magnetic field of $B_{0}=3\times 10^{18}$ G. Therefore, in our analysis we restrict the order of magnitude of $B_{0}$ to be at most three orders of magnitude larger than $B_{s}$. For the values of $\rho_{0}$ used in this work, our non-magnetized pure NS calculations give $\rho_{0}=3240.46~\text{MeV fm}^{-3}$ for $h=1$, $\rho_{0}=2008.27~\text{MeV fm}^{-3}$ for $h=1.5$, and $\rho_{0}=1459.81~\text{MeV fm}^{-3}$ for $h=2$. We see that all these values remain inside the range $\times 10^{-4}1×1$$\times 10^{4}~\text{MeV fm}^{-3}$. Thus, all values of $\rho_{0}$, $B_{s}$, and $B_{0}$ used throughout this study lie within the parameter ranges described above. The small value of $\rho_{0}$ is intentionally taken so that the chosen second term of the right-hand side of Eq. (LABEL:integrallongI1) remains applicable even in the small-order regime, without requiring concern over how small $\rho_{0}$ is.
@@
In Fig. 6, the gray region corresponds to the range $0\leq D<10^{-40}$, while the black region represents negative values of $D$. Note that $1\times 10^{-40}$ is the maximum threshold of numerical precision in our computation. We observe that no gray or black regions appear in the figure, indicating that the discriminant remains positive and always above $1\times 10^{-40}$ throughout all computed regimes. This shows that we can use the logarithmic function for the second term on the right-hand side of Eq. (LABEL:integrallongI1), instead of the inverse tangent function. Now Eq. (LABEL:integrallongI1) reads
In Appendix B, we show that for the isotropic configuration with the limit $\mathcal{B}\rightarrow 0$, $\left(e^{2\nu}\right)^{\prime}$ is reduces to the expression of $\left(e^{2\nu}\right)^{\prime}$ shown by Eq. (24) in Ref. Nojiri2024 .
@@
Furthermore, by substituting the expressions of $\alpha$, $\beta$, $\gamma$, and $z$ in Eq. (LABEL:alphabetagamma) into Eq. (76), we are left with
@@
A remark is stressed to be emphasized in order concerning the limit of vanishing magnetic field. In the magnetized configurations, the magnetic field modifies the geometric structure of the solution. As a result, the limit $\mathcal{B}\to 0$ is non-uniform: setting $\mathcal{B}=0$ directly at the level of intermediate quantities does not, in general, reproduce the non-magnetized configuration obtained by solving the full system.
As shown explicitly in A and B, the metric function $e^{2\nu}$ and its radial derivative $\left(e^{2\nu}\right)^{\prime}$ smoothly reduce to their non-magnetized counterparts when the limit $\mathcal{B}\to 0$ is taken at the level of the full metric functions. This point is crucial because the wormhole throat radius $r_{0}$ is determined from $\left(e^{2\nu}\right)^{\prime}$ evaluated at the surface $r=R_{s}$, where the energy density vanishes. At this stage, $\left(e^{2\nu}\right)^{\prime}$ contains the coefficients $\alpha$, $\beta$, and $\gamma$, which depend on the magnetic field and appear both in the numerator and the denominator, leading to the finite expression given in Eq. (77).
If the limit $\mathcal{B}\to 0$ is taken prematurely at the level of individual quantities, unphysical results may arise. For instance, the coefficient $e^{2\nu_{c}}$ appearing in the magnetized solution, shown by Eq. (LABEL:integrationconstantmagnetized), explicitly depends on the magnetic-field parameter $\gamma$ and would vanish identically if the limit were applied naively, causing the metric function $e^{2\nu}$ itself to collapse. This pathology is avoided once the limit is taken at the level of the full metric function, where the cancellation of magnetic-field–dependent terms occurs consistently, and the system is reduced to non-magnetized configuration.
This behavior shows that the magnetic field genuinely modifies the geometric structure of the system and that the non-magnetized configuration is recovered only when the limit is applied at the appropriate structural level. The magnetized and non-magnetized solutions therefore belong to distinct branches of the solution space, which justifies treating them separately in this work.
Note that in this work, regarding the influence of the magnetic field on the EOS, we follow the approach used in Refs. Mallick2014 ; Konno1999 ; Konno2000 ; Rizaldy2018b ; Rizaldy2019 , where the magnetic field contribution to the EOS is taken to be negligible. In particular, Refs. Konno1999 ; Konno2000 also employed a polytropic EOS in their analysis. The authors of Ref. Mallick2014 argued that the dominant magnetic effects arise from the additional stress and pressure terms, while the contribution from magnetization remains subdominant even for extremely strong magnetic fields, especially when such high central fields tend to drive the star toward instability. They used magnetic fields of the order of $10^{18}$ G in their analysis. The neglect of Landau quantization effects on the EOS has also been discussed in Ref. Sinha2013 .
In the numerical calculations of magnetized WH+NS configurations we consider two representative combinations of magnetic fields: (i) $B_{0}=3\times 10^{18}\,\mathrm{G}$ and $B_{s}=10^{15}\,\mathrm{G}$, and (ii) $B_{0}=3\times 10^{18}\,\mathrm{G}$ and $B_{s}=10^{16}\,\mathrm{G}$. Although the typical surface magnetic field of NSs lies in the range $10^{12}$$10^{15}\,\mathrm{G}$ as mentioned in Sec. 1, Ref. Gomes2019 reports that the limiting magnetic moment of NSs corresponds to a surface field of order $B_{s}\sim 10^{16}\,\mathrm{G}$. For this reason we also include $B_{s}=10^{16}\,\mathrm{G}$ among the cases examined in our numerical calculations. Configurations (i) and (ii) lie within the parameter ranges shown in Fig. 6. The anisotropy parameter is chosen to be the same as in the non-magnetized case, namely h=1, 1.5,h=1,\ 1.5, and $2$.
@@
Another important point is that observational constraints indicate that the typical radius of neutron stars lies in the range of approximately 1010-$11.5$ km ozel2016 , while the canonical neutron-star mass is $1.4\,M_{\odot}$ Pattersons2025 . Therefore, the very large radii in the range 20209090 km obtained in Fig. 7 are incompatible with current observational evidence, at least at the present stage.